# Back to the origins of brane-antibrane inflation

### v1

@@ Abstract
preprint: CERN-TH-2024-128
@@ 1 Introduction
$V(\varphi)=C_{0}\left(1-\frac{C_{1}}{\varphi^{n}}\right)\qquad\text{or}\qquad V(\varphi)=C_{2}\left(1-C_{3}\,e^{-m\varphi}\right),$ (1)
@@
Historically, following the seminal paper Dvali:1998pa , the first example of an inverse power-law potential in string cosmology has been actually proposed in the original formulation of brane-antibrane inflation which features $n=4$ Burgess:2001fx ; Dvali:2001fw . In that case the term proportional to $C_{0}$ is generated by the tension of an $\overline{\rm D3}$-brane, while the term proportional to $C_{1}$ corresponds to the Coulomb attraction between an $\overline{\rm D3}$-brane and a mobile D3-brane. However, it was soon realised that to get viable slow-roll inflation the distance between the brane and the antibrane should be larger than the typical size of the extra-dimensions. This problem was overcome in Kachru:2003sx where the authors placed the $\overline{\rm D3}$-brane at the tip of a highly warped throat. The authors of the same paper however pointed out that a stable inflationary trajectory requires to stabilise the modulus corresponding to the volume of the extra dimensions. In turn, the dynamics responsible to fix the volume mode can induce inflaton-dependent contributions to the scalar potential which in general ruin its flatness. This is the famous $\eta$-problem that plagues several inflationary models in supergravity and string theory. In particular, after including these new $\varphi$-dependent contributions, the resulting potential does not feature anymore an inverse power-law behaviour with $n=4$. A more refined analysis Baumann:2007ah showed that concrete models of brane-antibrane inflation can have enough tuning freedom to realise inflection point inflation. This is however a scenario that is very different from the original one proposed in Burgess:2001fx ; Dvali:2001fw and, above all, it is disfavoured by observations. Nevertheless, when comparing general models with observations, it has remained a practice in the literature to consider inverse power-law potentials still under the name of brane-antibrane inflation Planck:2018jri ; Martin:2024qnn even in the absence of a string realisation of these models once moduli stabilisation is taken into account.
@@ 2 A historical perspective on brane-antibrane inflation
$C_{0}\equiv 2T_{3}\,\mathcal{V}^{2/3}\,e^{-8\pi K/(3g_{s}M)}\qquad\text{and}\qquad D_{0}\equiv\frac{27}{32\pi^{2}T_{3}}\,\mathcal{V}^{2/3}\,e^{-8\pi K/(3g_{s}M)}\,,$ (7)
@@
$V_{\rm tot}=V_{\rm np}+V_{\rm inf}=\frac{1}{\left(T+\bar{T}-\gamma r^{2}\right)^{2}}\left[U_{\rm np}(T+\bar{T})+U_{\rm inf}(r)\right].$ (17)
@@
where $V_{0}(\varphi)$ is given by (14) with fixed $\mathcal{V}$. It is then straightforward to realise that the Planck-suppressed 6D operator $V_{0}(\varphi)\varphi^{2}$ induces a large contribution to the slow-roll parameter $\eta$ of order $\Delta\eta=2/3$ which ruins inflation. Ref. Kachru:2003sx argued however that the prefactor of non-perturbative corrections to $W$ is in general a function of $\varphi$, and so one should have more precisely that $U_{\rm np}=U_{\rm np}(T+\bar{T},\varphi)=U_{\rm np}(T+\bar{T})+U_{\rm sub}(T+\bar{T},\varphi)$ where we included the $\varphi$-dependence in a subleading term. Thus the expansion of the scalar potential should contain an additional term of the form:
$V_{\rm inf}\simeq V_{0}(\varphi)\left(1+\frac{1}{3}\,\frac{\varphi^{2}}{M_{p}^{2}}+P(\varphi)\right)\qquad\text{where}\qquad P(\varphi)\equiv\frac{U_{\rm sub}(\langle(T+\bar{T})\rangle,\varphi)}{U_{\rm inf}(\varphi)}\,.$ (21)
@@ 3 Late-time moduli stabilisation
    $\tau=\mathcal{V}^{2/3}\qquad\to\qquad\tau=\mathcal{V}^{2/3}+c_{0}\,,$ (23)
@@
    $\tau\quad\to\quad\tau-\alpha\ln\tau\qquad\Rightarrow\qquad K=-3\ln\tau\quad\to\quad K=-3\ln\left(\tau-\alpha\ln\tau\right),$ (24)
@@
    $K_{\mathcal{O}(g_{s}^{n}\alpha^{\prime 3})}=-2\ln\left\{1+\frac{\xi}{2(g_{s}\tau)^{3/2}}\left[1+g_{s}^{2}\left(c_{2}\left(1-\frac{3T_{7}}{2}\ln\tau\right)+c_{3}\right)\right]\right\},$ (27)
@@
    $\xi=-\frac{\zeta(3)}{2(2\pi)^{3}}\left(\chi(\rm{CY})+2\int_{\rm CY}D_{\rm O7}^{3}\right)\qquad\text{and}\qquad c_{2}=\frac{2\zeta(2)}{\zeta(3)}\,,$ (28)
@@
$\frac{V}{3W_{0}^{2}}=e^{K}\left(\frac{1}{3}K^{T\overline{T}}K_{T}K_{\overline{T}}-1\right)\simeq\frac{\alpha}{\tau^{4}}-\frac{\xi\sqrt{g_{s}}}{4c\tau^{9/2}}\left(\ln\tau-\frac{c}{g_{s}^{2}}\right)\,.$ (33)
@@
$\tau_{\rm min}=\lambda_{0}\,e^{\frac{c}{g_{s}^{2}}}\qquad\tau_{\rm max}=\lambda_{-1}\,e^{\frac{c}{g_{s}^{2}}}\qquad\text{with}\qquad\lambda_{k}\equiv e^{\frac{2}{9}}\,e^{-2\mathcal{W}_{k}\left(-\theta/e\right)}\,,$ (34)
@@
$\frac{81\theta_{*}}{8\left(e^{1+\mathcal{W}_{0}(-\theta_{*}/e)}+9\theta_{*}\right)}>1\qquad\Rightarrow\qquad\theta_{*}\simeq 0.993<\theta<1\,.$ (36)
@@
$\theta\simeq 1\qquad\Leftrightarrow\qquad\alpha\simeq 4.78\sqrt{g_{s}}\,e^{-\frac{0.02}{g_{s}^{2}}}\simeq 0.22\qquad\text{for}\qquad g_{s}\simeq 0.1\,,$ (37)
which for example would be exactly in the right ballpark if $\alpha$ were given in terms of the 1-loop $\beta$-function coefficient of an $SU(2)$ theory as $\alpha=\beta_{0}/(8\pi)=3/(4\pi)\simeq 0.24$. We therefore conclude that late-time dS minima should naturally exist in the string landscape for appropriate values of the flux quanta which determine $g_{s}$. The behaviour of the late-time scalar potential for different values of $\theta$ is presented in Fig. 1. Let us stress that (34) receives subleading corrections but in App. B we show that subdominant contributions do not modify our results qualitatively - rather they make our scenario more robust.
@@ 4 Early-time moduli stabilisation: brane-antibrane inflation
To do so, we need to consider three fields: the Kähler modulus $\tau\equiv(T+\bar{T})$, the inflaton $r$ that controls the distance of the D3-brane from the $\overline{\rm D3}$-brane at the tip of the throat, and a nilpotent superfield $X$ which describes the dynamics of the $\overline{\rm D3}$-brane and realises supersymmetry non-linearly. In fact, $X$ is subject to the constraint $X^{2}=0$ that projects out all bosonic degrees of freedom, leaving just the Goldstino in the low-energy EFT. [^9] Ref. Burgess:2022nbx considered instead the case where the inflaton $r$ is the scalar component of a chiral superfield $\Phi$ subject to the constraint $\overline{\mathcal{D}}(X\bar{\Phi})=0$ which projects out the fermionic and auxiliary fields. In our treatment we do not need to include the inflaton in the sector which realises supersymmetry non-linearly.
@@
It is this quantity, $\sigma$, and not $\tau$, that is stabilised perturbatively during inflation, guaranteeing the absence of the $\eta$-problem. Note that after the end of inflation $r\to 0$ and $\sigma\to\tau$. The 4D supergravity EFT can be expanded in powers of $X$ as follows Burgess:2021obw ; Burgess:2022nbx ; DallAgata:2015zxp :
@@
$e^{-K}\,V=\left.\left[K^{\overline{X}X}\,W_{X}^{2}+W_{0}W_{X}\left(K^{\overline{X}A}K_{A}+K^{\overline{A}X}K_{\overline{A}}\right)+W_{0}^{2}\left(K^{\overline{A}B}K_{\overline{A}}K_{B}-3\right)\right]\right|_{X=0},$ (40)
@@
$V=\frac{1}{U}\left[\left(f^{\prime}W_{X}-3g^{\prime}W_{0}\right)^{2}-f^{\prime\prime}\left(fW_{X}^{2}-6gW_{X}W_{0}-9hW_{0}^{2}\right)\right],$ (41)
@@
$W_{X}(r)=e^{-2\rho}\sqrt{\frac{3}{4\pi}\left(1-\frac{\mathcal{D}_{0}}{\left(rM_{\scriptscriptstyle KK}\right)^{4}}\right)}\qquad\text{with}\qquad\rho\equiv\frac{2\pi K}{3g_{s}M}\,,$ (44)
the potential (43) reproduces the inflationary potential (11). However $\sigma$ would be an unstable runaway direction at tree-level. We need therefore to add quantum corrections to fix $\sigma$. In what follows we shall analyse two separate cases: the first with a linear dependence of the Kähler potential on the nilpotent superfield $X$, i.e. $g(\sigma)\neq 0$, while the second without a contribution in $K$ which is linear in $X$, i.e. $g(\sigma)=0$.
@@ 4.1 Case with linear contribution
$\sigma_{\rm min}=3g_{s}\frac{W_{0}}{W_{X}}\simeq 2\sqrt{3\pi}\,g_{s}W_{0}\,e^{2\rho}\gg 1\qquad\text{for}\qquad e^{2\rho}\gg 1\,.$ (46)
@@
$f(\sigma)=\sigma+\frac{\xi}{3g_{s}^{3/2}\sqrt{\sigma}}\,,\qquad g(\sigma)=g_{s}\ln\sigma\,,\qquad h(\sigma)=1\,.$ (47)
@@
$C_{0}\equiv\frac{9\xi}{192\sqrt{2}(3\pi)^{9/4}g_{s}^{6}}\frac{e^{-9\rho}}{W_{0}^{5/2}}\qquad\text{and}\qquad C_{1}\equiv\frac{9\mathcal{D}_{0}T_{3}^{2}}{4M_{\scriptscriptstyle KK}^{4}}\,.$ (51)
This potential scales as the original brane-antibrane potential (43) but with the difference that now $\sigma$ is stabilised perturbatively at $\sigma=\sigma_{\rm min}$. This has two very important implications: ($i$) inflaton can occur with a stable volume direction; ($ii$) this model is not plagued by any $\eta$-problem since perturbative corrections to the effective action stabilise $\sigma=\tau-\gamma r^{2}$, instead of just $\tau$.
@@
Figure 3: The blue curve is the inflationary potential for the case with linear contribution which features a dS minimum with $V_{\rm min}^{\rm inf}\sim 10^{-15}$, while the yellow curve is the late-time potential with a smaller vacuum energy at $V_{\rm min}\sim 10^{-16}$ and a minimum at larger field values. The parameters used to generate this plot are $\xi=g_{s}=0.1$, $W_{0}=1$, $W_{X}=10^{-3}$ and $T_{7}=2\pi$, while $\alpha$ has been fixed as in (35) with $\theta=0.994$.

### v2

@@ Abstract
pacs:
PACS-keydiscribing text of that key and PACS-keydiscribing text of that key
@@ 1 Introduction
$V(\varphi)=C_{0}\left(1-\frac{C_{1}}{\varphi^{n}}\right)\quad\text{or}\quad V(\varphi)=C_{2}\left(1-C_{3}\,e^{-m\varphi}\right),$ (1)
@@
Historically, following the seminal paper Dvali:1998pa , the first example of an inverse power-law potential in string cosmology has been actually proposed in the original formulation of brane-antibrane inflation which features $n=4$ Burgess:2001fx ; Dvali:2001fw . In that case the term proportional to $C_{0}$ is generated by the tension of an $\overline{\rm D3}$-brane, while the term proportional to $C_{1}$ corresponds to the Coulomb attraction between an $\overline{\rm D3}$-brane and a mobile D3-brane.
This scenario is very appealing since it not only had a concrete string theory candidate for inflation (the brane separation), it also had a very stringy way to end inflation by means of an open string mode that becomes tachyonic at a critical separation Burgess:2001fx therefore realising a version of hybrid inflation. However, it was soon realised Burgess:2001fx ; Dvali:2001fw that to get viable slow-roll inflation the distance between the brane and the antibrane should be larger than the typical size of the extra-dimensions. This problem was overcome in Kachru:2003sx where the authors placed the $\overline{\rm D3}$-brane at the tip of a highly warped throat. The authors of the same paper however pointed out that a stable inflationary trajectory requires to stabilise the modulus corresponding to the volume of the extra dimensions. In turn, the dynamics responsible to fix the volume mode can induce inflaton-dependent contributions to the scalar potential which in general ruin its flatness.
This is the famous $\eta$-problem that plagues several inflationary models in supergravity and string theory. In particular, after including these new $\varphi$-dependent contributions, the resulting potential does not feature anymore an inverse power-law behaviour with $n=4$. A more refined analysis Baumann:2007ah showed that concrete models of brane-antibrane inflation can have enough tuning freedom to realise inflection point inflation. This is however a scenario that is very different from the original one proposed in Burgess:2001fx ; Dvali:2001fw and, above all, it is disfavoured by observations. Nevertheless, when comparing general models with observations, it has remained a practice in the literature to consider inverse power-law potentials still under the name of brane-antibrane inflation Planck:2018jri ; Martin:2024qnn even in the absence of a string realisation of these models once moduli stabilisation is taken into account.
@@ 2 A historical perspective on brane-antibrane inflation
$C_{0}\equiv 2T_{3}\,\mathcal{V}^{2/3}\,e^{-8\pi K/(3g_{s}M)}\,,$ (7)
and:
$D_{0}\equiv\frac{27}{32\pi^{2}T_{3}}\,\mathcal{V}^{2/3}\,e^{-8\pi K/(3g_{s}M)}\,,$ (8)
@@
$V_{\rm tot}=V_{\rm np}+V_{\rm inf}=\frac{1}{\left(T+\bar{T}-\gamma r^{2}\right)^{2}}\left[U_{\rm np}(T+\bar{T})+U_{\rm inf}(r)\right]$ (18)
@@
where $V_{0}(\varphi)$ should be identified with the leading inflationary potential given by (15) with fixed $\mathcal{V}$. It is then straightforward to realise that the Planck-suppressed 6D operator $V_{0}(\varphi)\varphi^{2}$ induces a large contribution to the slow-roll parameter $\eta$ of order $\Delta\eta=2/3$ which ruins inflation. Ref. Kachru:2003sx argued however that the prefactor of non-perturbative corrections to $W$ is in general a function of $\varphi$, and so one should have more precisely that $U_{\rm np}=U_{\rm np}(T+\bar{T},\varphi)=U_{\rm np}(T+\bar{T})+U_{\rm sub}(T+\bar{T},\varphi)$ where we included the $\varphi$-dependence in a subleading term. Thus the expansion of the scalar potential should contain an additional term of the form:
$V_{\rm inf}\simeq V_{0}(\varphi)\left(1+\frac{1}{3}\,\frac{\varphi^{2}}{M_{p}^{2}}+P(\varphi)\right)$ (22)
where:
$P(\varphi)\equiv\frac{U_{\rm sub}(\langle(T+\bar{T})\rangle,\varphi)}{U_{\rm inf}(\varphi)}\,.$ (23)
@@ 3 Late-time moduli stabilisation
    $\tau=\mathcal{V}^{2/3}\quad\to\quad\tau=\mathcal{V}^{2/3}+c_{0}\,,$ (25)
@@
    $\tau\quad\to\quad\tau-\alpha\ln\tau$ (26)
    that implies:
    $K=-3\ln\tau\quad\to\quad K=-3\ln\left(\tau-\alpha\ln\tau\right),$ (27)
@@
    $K_{\mathcal{O}(g_{s}^{n}\alpha^{\prime 3})}$ $=$ $-2\ln\left\{1+\frac{\xi}{2(g_{s}\tau)^{3/2}}\right.$
    $\left.\left[1+g_{s}^{2}\left(c_{2}\left(1-\frac{3T_{7}}{2}\ln\tau\right)+c_{3}\right)\right]\right\},$
@@
    $\xi=-\frac{\zeta(3)}{2(2\pi)^{3}}\left(\chi(\rm{CY})+2\int_{\rm CY}D_{\rm O7}^{3}\right)\,\,\text{and}\,\,c_{2}=\frac{2\zeta(2)}{\zeta(3)},$ (31)
@@
$\frac{V}{3W_{0}^{2}}$ $=$ $e^{K}\left(\frac{1}{3}K^{T\overline{T}}K_{T}K_{\overline{T}}-1\right)$ (36)
$\simeq$ $\frac{\alpha}{\tau^{4}}-\frac{\xi\sqrt{g_{s}}}{4c\tau^{9/2}}\left(\ln\tau-\frac{c}{g_{s}^{2}}\right).$
@@
$\tau_{\rm min}=\lambda_{0}\,e^{\frac{c}{g_{s}^{2}}}\qquad\text{and}\qquad\tau_{\rm max}=\lambda_{-1}\,e^{\frac{c}{g_{s}^{2}}}\,,$ (37)
with:
$\lambda_{k}\equiv e^{\frac{2}{9}}\,e^{-2\mathcal{W}_{k}\left(-\theta/e\right)}\,,$ (38)
@@
$\frac{81\theta_{*}}{8\left(e^{1+\mathcal{W}_{0}(-\theta_{*}/e)}+9\theta_{*}\right)}>1\quad\Rightarrow\quad\theta_{*}\simeq 0.993<\theta<1\,.$ (40)
@@
$\theta\simeq 1\quad\Leftrightarrow\quad\alpha\simeq 4.78\sqrt{g_{s}}\,e^{-\frac{0.02}{g_{s}^{2}}}\simeq 0.22\quad\text{for}\quad g_{s}\simeq 0.1\,,$ (41)
which for example would be exactly in the right ballpark if $\alpha$ were given in terms of the 1-loop $\beta$-function coefficient of an $SU(2)$ theory as $\alpha=\beta_{0}/(8\pi)=3/(4\pi)\simeq 0.24$. Note that similar values of $\alpha$ can be obtained for different values of $\xi$ and $g_{s}$ (for example $\xi=0.1$ would require $g_{s}\sim 0.3$). We therefore conclude that late-time dS minima should naturally exist in the string landscape for appropriate values of the flux quanta which determine $g_{s}$. The behaviour of the late-time scalar potential for different values of $\theta$ is presented in Fig. 1. Let us stress that (37) receives subleading corrections but in App. B we show that subdominant contributions do not modify our results qualitatively - rather they make our scenario more robust.
@@ 4 Early-time moduli stabilisation: brane-antibrane inflation
To do so, we need to consider three fields: the Kähler modulus $\tau\equiv(T+\bar{T})$, the inflaton $r$ that controls the distance of the D3-brane from the $\overline{\rm D3}$-brane at the tip of the throat, and a nilpotent superfield $X$ which describes the dynamics of the $\overline{\rm D3}$-brane and realises supersymmetry non-linearly. [^9] See Antoniadis:2024hvw for a recent review and Guleryuz:2022hiv for recent studies on non-linear supergravity and cosmology. In fact, $X$ is subject to the constraint $X^{2}=0$ that projects out all bosonic degrees of freedom, leaving just the Goldstino in the low-energy EFT. [^10] Ref. Burgess:2022nbx considered instead the case where the inflaton $r$ is the scalar component of a chiral superfield $\Phi$ subject to the constraint $\overline{\mathcal{D}}(X\bar{\Phi})=0$ which projects out the fermionic and auxiliary fields. In our treatment (see also Aparicio:2015psl ) we do not need to include the inflaton in the sector which realises supersymmetry non-linearly. In fact, due to the geometrical separation between the D3- and the $\overline{\rm D3}$-brane which realises supersymmetry non-linearly, we expect the inflaton to be the component of a standard chiral superfield which realises supersymmetry linearly.
@@
It is this quantity, $\sigma$, and not $\tau$, that is stabilised perturbatively during inflation, guaranteeing the absence of the $\eta$-problem. Note that after the end of inflation $r\to 0$ and $\sigma\to\tau$. The 4D supergravity EFT can be expanded in powers of $X$ as follows Burgess:2022nbx ; Burgess:2021obw ; DallAgata:2015zxp ; McDonough:2016der ; Kallosh:2017wnt :
@@
$e^{-K}\,V$ $=$ $\left[K^{\overline{X}X}\,W_{X}^{2}+W_{0}W_{X}\left(K^{\overline{X}A}K_{A}+K^{\overline{A}X}K_{\overline{A}}\right)\right.$ (44)
$+$ $\left.\left.W_{0}^{2}\left(K^{\overline{A}B}K_{\overline{A}}K_{B}-3\right)\right]\right|_{X=0},$
@@
$V$ $=$ $\frac{1}{U}\left[\left(f^{\prime}W_{X}-3g^{\prime}W_{0}\right)^{2}\right.$ (45)
-- $\left.f^{\prime\prime}\left(fW_{X}^{2}-6gW_{X}W_{0}-9hW_{0}^{2}\right)\right],$
@@
$W_{X}(r)=e^{-2\rho}\sqrt{\frac{3}{4\pi}\left(1-\frac{\mathcal{D}_{0}}{\left(rM_{\scriptscriptstyle KK}\right)^{4}}\right)}\,,$ (48)
with:
$\rho\equiv\frac{2\pi K}{3g_{s}M}\,,$ (49)
the potential (47) reproduces the inflationary potential (12). Let us stress that the form of $W_{X}$ is inferred by requiring that the scalar potential term of the DBI $\overline{\rm D3}$-brane action should also be expressible as a standard supergravity F-term potential. However $\sigma$ would be an unstable runaway direction at tree-level. We need therefore to add quantum corrections to fix $\sigma$. In what follows we shall analyse two separate cases: the first with a linear dependence of the Kähler potential on the nilpotent superfield $X$, i.e. $g(\sigma)\neq 0$, while the second without a contribution in $K$ which is linear in $X$, i.e. $g(\sigma)=0$.
@@ 4.1 Case with linear contribution
$\sigma_{\rm min}=3g_{s}\frac{W_{0}}{W_{X}}\simeq 2\sqrt{3\pi}\,g_{s}W_{0}\,e^{2\rho}\gg 1\quad\text{for}\quad e^{2\rho}\gg 1\,.$ (51)
@@
$f(\sigma)=\sigma+\frac{\xi}{3g_{s}^{3/2}\sqrt{\sigma}}\,,\quad g(\sigma)=g_{s}\ln\sigma\,,\quad h(\sigma)=1\,.$ (52)
@@
$C_{0}\equiv\frac{9\xi}{192\sqrt{2}(3\pi)^{9/4}g_{s}^{6}}\frac{e^{-9\rho}}{W_{0}^{5/2}}\quad\text{and}\quad C_{1}\equiv\frac{9\mathcal{D}_{0}T_{3}^{2}}{4M_{\scriptscriptstyle KK}^{4}}\,.$ (56)
This potential scales as the original brane-antibrane potential (47) but with the difference that now $\sigma$ is stabilised perturbatively at $\sigma=\sigma_{\rm min}$. This has two very important implications: ($i$) inflation can occur with a stable volume direction; ($ii$) this model is not plagued by any $\eta$-problem since perturbative corrections to the effective action stabilise $\sigma=\tau-\gamma r^{2}$, instead of just $\tau$.
@@
Figure 3: The blue curve is the inflationary potential for the case with linear contribution which features a dS minimum with $V_{\rm min}^{\rm inf}\simeq 2.5\times 10^{-15}$, while the yellow curve is the late-time potential with a smaller vacuum energy at $V_{\rm min}\simeq 8\times 10^{-16}$ and a minimum at larger field values. The parameters used to generate this plot are $\xi=g_{s}=0.1$, $W_{0}=1$, $W_{X}=10^{-3}$ and $T_{7}=2\pi$, while $\alpha$ has been fixed as in (39) with $\theta=0.994$.